OpenStudy (anonymous):

Which expression shows how 5 • 37 can be rewritten using the distributive property? A. 5 • 30 + 5 • 7 B. 5 • 30 + 5 • 37 C. 5 • 30 + 7 D. 15 • 30 + 15 • 7

jigglypuff314 (jigglypuff314):

(5) * (37) = (5) * (30 + 7) then distribute (Don't confuse with C, parenthesis matter) = (5)*(30) + (5)*(7) <---